Phonebill

logfile analyser Version 2.0

27 September 1993

by Raymond Penners

1 Introduction

‘Phonebill’ is, simply said, a logfile analyser. What it basically does is scan the logfile(s) generated by a terminal program or a mailer, extract all information about calls you have made by using your modem, and store this in its own (short) format. Features:

• Requires Kickstart 2.04 or higher.

• Supports new 3.0 features (new look menus, ...).

• Nice gadtools compliant GUI.

• User-definable callrates, supporting rate exceptions for certain days and dates. Handles fees with up to 4 decimals (e.g. $1,2344 dollars per 30 seconds) for extra accuracy.

• Supports logfiles generated by:

• Section 7.2 [AmigaUUCP Timelog], page 11

• Section 7.4 [JrComm], page 12

• Section 7.5 [MagiCall], page 13

• Section 7.7 [NComm], page 14

• Section 7.1 [TrapDoor], page 11

• Term (Section 7.6 [Term action-log], page 13 and Section 7.8 [Term call-log], page 14)

• Section 7.3 [Terminus], page 12

• Automatical logfile truncating.

• Generates miscellaneous reports: statistics, total costs, etc.

• Context-sensitive online help.

2 Background

‘Phonebill’ was created for one simple reason: I couldn’t find a good logfile analyser. Sure, there are hundreds of logfile analysers that sometimes even calculate the costs correctly, but all these programs are just simple, very limited, uncomfortable, mostly CLI-based programs, that don’t offer the features I need.

3 Installation

Just drag the ‘Phonebill’ drawer over to the drawer where you would like to copy it to. That’s all there is to it!

3.1 Upgrading from version 1.0

Please note that all the configuration files have changed. This means that you can’t use your V1.0 configuration files anymore. You may delete all phonebill files found in the S: drawer, except for the ‘S:Phonebill.Log’ file. I am sorry for the inconvenience. This will not happen again.

3.2 Upgrading from version 1.07 - 1.10

Starting from version 2.0 there is a complete new rate editor. Due to the fact that the old rates are totally different from the new rates ‘Phonebill’ can’t read the v1.x rates anymore. However, the program still loads the host configuration, resetting the hosts’ rate. You can reinstall the missing information in the host manager, or you could delete the ‘ENV:Phonebill/HostRate.Prefs’ file so that you will be prompted for the missing information automatically when ‘Phonebill’ needs it.

4 Workbench and Shell

‘Phonebill’ can be started from the Shell or the workbench. When started from the Shell, ‘Phonebill’ accepts commandline arguments. To view all known arguments, supply a "?" as only argument to the program, for example:

        1.SPHINX-O-HD:Comm/Misc> Phonebill ?

So far only one argument is accepted:

• Process/S Phonebill will process the logfiles, and return control to the Shell when done.

5 Menus

The following sections describe the menu items available in the ‘Phonebill’ main panel.

5.1 Project

Save configuration

Saves the program configuration.

Iconify Closes the window and drops an icon labeled ‘Phonebill’ on the Workbench window. Double-clicking this icon will cause ‘Phonebill’ to wake up and to return the state it was in before ‘Phonebill’ was iconified.

About Shows some additional information on the program.

Quit Terminates the program. If the configuration has been changed, you will be prompted to confirm your decision.

5.2 Edit

Preferences

See Section 6.2 [Preferences panel], page 6.

Log manager

See Section 6.3 [Log manager panel], page 7.

Host manager

See Section 6.5 [Host manager panel], page 8.

Default Resets the program configuration to its defaults.

Last saved

Loads the last saved configuration.

6 Control panels

6.1 Main panel

Here it all begins. This panel provides gadgets and Chapter 5 [Menus], page 5 which give you control over the main functions of ‘Phonebill’.

Process Logs

Scans and/or truncates all logfiles listed in the log manager, See Section 6.3 [Log manager panel], page 7. If new calls were found ‘Phonebill’ will add them to its own binary logfile.

Preferences

See Section 6.2 [Preferences panel], page 6.

About Shows some additional information on the program.

From, To The dates entered here specify the scanrange. Only calls made within the period <from date> - <to date> will be used to generate the report. Please note that when you are using a localised workbench you’ll have to enter localised dates. For example: "monday" has to be "maandag" when using a Dutch workbench.

Statistics, Phonebill, List calls

If you select on of these ‘Phonebill’ will start to generate a report. This report will be shown to you using the text viewer entered in the preferences.

Maintenance

See Section 6.8 [Maintenance panel], page 10.

6.2 Preferences panel

TextViewer

The textviewer to be used for displaying the report generated by the ‘Phonebill’, ‘List Calls’ and ‘Statistics’ options.

Report file

The filename of the report to be generated.

Log file The filename of the binary log file generated by phonebill.

CON: Window

Specifies the console window which will be used for miscellaneous output.

Carrier detect time

It always takes some time before your modem has detected a carrier. The connect time in the logfiles generated by your terminal program are therefore incorrect, because the modem has been online trying to get a carrier for a couple of seconds. For example, it takes my 14K4 (V.32bis) approximately 15 seconds to connect.

Delete report after viewing

If this option is checked, the report will be deleted after it has been viewed.

Save <from> and <to> date on exit

If this option is checked, the <from> and <to> date will be saved automatically when you exit the program.

Log manager

See Section 6.3 [Log manager panel], page 7.

Host manager

See Section 6.5 [Host manager panel], page 8.

Save Saves currents preferences, and closes the preferences panel. The from/to dates from the main window are also saved.

Use Closes the preferences panel, without saving the preferences.

6.3 Log manager panel

In the Log Manager panel you can add, remove and edit logfiles and their options, for use with the ‘Phonebill’ program.

Add Adds a new logfile to the list.

Edit Lets you edit the settings of the selected logfile, See Section 6.4 [Log settings panel], page 7.

Remove Removes the selected logfile.

Save Saves current log settings, and closes the log manager panel.

Use Closes the log manager panel, without saving the log settings.

6.4 Log settings panel

Logfile The filename of the logfile to be used.

Log format

Specifies the type of the logfile. See Chapter 7 [Supported logfiles], page 11.

Max. logsize

The maximum size (in KB) of the logfile. If its size exceeds this value, the file will be truncated to 50% of the maximum size. This only works if the truncate option is activated.

Scan log for phonecalls

If this option is enabled, the logfile will be scanned for calls.

Truncate log if maximumsize is reached
          See Max. logsize.
Use archive bit to check if file changed

If enabled, Phonebill will set the archive bit of the logfile after processing. When it processes this file again it recognizes wether or not the file has changed. This will speed-up the processing.

6.5 Host manager panel

In the host manager panel you see a list of all ‘known’ hosts. You can add,remove and edit hosts.

Add Adds a new hosts to the list.

Edit Lets you edit the settings of the selected host, See Section 6.6 [Host settings panel], page 8.

Remove Removes the selected host.

Save Saves current host settings, and closes the host manager panel.

Use Closes the host manager panel, without saving the host settings.

6.6 Host settings panel

Name, Number

The name and/or telephonenumber of the host. You must at least fill in one of them.

Rate list This list contains all installed rates. To select a rate for the host, simply choose on of the rates in the list.

Add Adds a new rate to the list.

Copy The selected rate will be duplicated.

Edit Lets you edit the selected rate, See Section 6.7 [Rate panel], page 9.

Remove Removes the selected rate.

Ok,Cancel

Use or drop the current host settings.

6.7 Rate panel

The information to be specified in this control panel will be used to calculate the amount of money to be paid for a call.

Rate name Specifies the name of the rate.

Days and dates

This list contains the default rate settings and exceptions for certain dates and days of the week. Each line displays the type of the entry and a comment (seperated by the ‘>>’ character). The following types are available:

Day(s) Settings for certain days of the week.

          17 Jul (example)

Settings for a specific date.

If there is no special type available for an entry, it’s probably the default settings you are dealing with. These settings are used whenever ‘Phonebill’ can’t find an entry for the current day.

For each entry in this list there is at least one associated starting time available which defines when the associated rate settings are to be used. You will find the time settings in the list titled Time. To add a new time use the Add time button, to edit an existing entry, use the Edit time button. To remove an entry, press the Remove time button.

Add date Will invoke a control panel to create a new rate entry to be used on a specific date.

Use the sliders to select the day the settings will be valid for.

Add day(s)

Will create a new rate entry referring to one or more days of the week. Use the buttons of the control panel to select the days the current settings will be valid for.

Copy Duplicates the selected entry and appends it to the list.

Edit Lets you modify the selected entry.

Remove Removes the selected entry from the list.

Cost/unit

The amount of money to be paid for each single time unit when online. This fee may contain a number followed by a "," or a ".", followed by up to 4 decimals. For example: if you have to pay 3 dollars and 14,15 cents per unit enter "3,1415".

Seconds/unit

This is where you enter how many seconds each time unit lasts.

There are two different groups of the two entries listed above: one for the first unit and one for the following units. So, if you only pay for the call you make but not for the time you spend online, just enter the fee in the first group and set the second group to zero.

6.8 Maintenance panel

In the maintenance panel you can edit the binary logfile. Only calls that fall within the specified scanrange in the main panel will be shown in the list.

Remove Removes the selected call from the list.

Remove all

Removes all calls from the list.

Save Saves the binary logfile.

Cancel Closes the maintenance panel without saving any changes.

7 Supported logfiles

This chapter contains an example of each of the supported logfiles. If you experience logscan problems, or if you don’t know what type of logfile you have compare it to on of the following.

7.1 TrapDoor

+ 11-Jul-93 12:05:10  TrapDoor 1.83 opened
+ 11-Jul-93 12:05:14  Calling 046-581998
~ 11-Jul-93 12:05:25  BUSY
+ 11-Jul-93 12:05:32  TrapDoor 1.83 closed
+ 11-Jul-93 12:05:55  TrapDoor 1.83 opened
+ 11-Jul-93 12:05:59  Calling 05430-24097
~ 11-Jul-93 12:06:29  CONNECT 14400/ARQ
: 11-Jul-93 12:06:34   Name: The Amiga Workbench (2:283/410.0)
: 11-Jul-93 12:06:34    AKA: 14:105/201.0 39:157/1.0 39:157/100.0
: 11-Jul-93 12:06:34  Sysop: Robert Udo
: 11-Jul-93 12:06:34  Place: Winterswijk - Gelderland (31-5430-24097)
: 11-Jul-93 12:06:34  Flags: CM
: 11-Jul-93 12:06:34  Using: InterMail 2.22
: 11-Jul-93 12:06:34  Offer: Emsi FReqs ZedZap
| 11-Jul-93 12:06:37  Receiving 000c043c.pkt (4410 bytes)
| 11-Jul-93 12:06:40  Took 0:02, Cps: 1575, Efficiency: 109%
= 11-Jul-93 12:07:10  Session connect time 0:40
+ 11-Jul-93 12:07:14  TrapDoor 1.83 closed
+ 11-Jul-93 12:15:45  PhoneBill LogScan

7.2 AmigaUUCP Timelog

01-May-93 12:13  0:09 in=0        out=0        cbmger
01-May-93 12:15  0:19 in=1496     out=0        cbmger
02-May-93 00:05  0:18 in=0        out=0        cbmger
02-May-93 00:35  0:08 in=0        out=0        cbmger
02-May-93 00:38  0:43 in=0        out=0        cbmger
02-May-93 01:14  0:30 in=0        out=722      cbmger
02-May-93 03:01  0:22 in=790      out=0        cbmger
07-May-93 15:03  2:42 in=122857   out=0        cbmger
07-May-93 16:18  0:28 in=5014     out=370      cbmger
07-May-93 16:21  0:23 in=3847     out=0        cbmger
07-May-93 16:36  0:22 in=0        out=1096     cbmger
07-May-93 16:41  1:33 in=82352    out=0        cbmger
07-May-93 17:34  3:15 in=111215   out=0        cbmger

08-May-93 03:15  2:14 in=128822   out=0        cbmger
11-May-93 19:59 12:24 in=707799   out=0        cbmger

7.3 Terminus

21-07-93 16:58:14 Session start
21-07-93 16:59:59 Connection at 14400bps to: SomeBBS  5485321
21-07-93 17:01:00 Carrier lost
21-07-93 20:49:59 Hangup modem
21-07-93 20:49:59 Session end
21-07-93 21:35:20 Session start
21-07-93 21:35:58 Connection at 14400bps to: Suncity  5489171
21-07-93 21:36:58 ZMODEM d/l opts: Binary - Overwrite
21-07-93 21:36:58 ZMODEM download: hd1:comm/xfer/pbill999.lha
21-07-93 21:37:02 Completed transfer - size: 530, time: 00:00:04, cps: 372
21-07-93 21:37:35 Hangup modem
21-07-93 21:37:35 Carrier lost
21-07-93 22:48:35 Phonebill LogScan

7.4 JrComm

07/20/93 01:34:05 - JR-Comm session started.
07/20/93 01:36:28 - Amigoline HQ:
07/20/93 01:39:49 - Amigoline HQ:
07/20/93 01:58:06 - Z d/l: 0/199 -          978 - ideeen.doc
07/20/93 01:58:19 - Z d/l: 0/215 -         2111 - inleidin.doc
07/20/93 02:03:20 - Z d/l: 0/232 -        64319 - vcheck626.lha
07/20/93 02:24:49 - JR-Comm session finished.
07/20/93 02:29:38 - JR-Comm session started.
07/20/93 02:29:59 - Amigoline HQ:
07/20/93 02:38:23 - JR-Comm session finished.
07/20/93 03:55:44 - JR-Comm session started.
07/20/93 03:56:13 - megatron bbs:
07/20/93 04:07:05 - Z u/l: 0/231 -        23654 - Dimp.lha
07/20/93 04:20:43 - Z d/l: 1/219 -        91783 - Kinski.lha
07/20/93 04:21:28 - Z d/l: 0/225 -         9441 - Nagel2.lha
07/20/93 04:22:10 - Z d/l: 0/226 -         8966 - Nagel3.lha
07/20/93 04:23:05 - Z d/l: 0/228 -        11995 - Sexyrobot.lha
07/20/93 04:28:09 - Z d/l: 0/231 -        69861 - Corinne.lha
07/20/93 04:28:33 - JR-Comm session finished.
07/20/93 16:01:50 - JR-Comm session started.
07/20/93 16:03:14 - Amigoline HQ:
07/20/93 16:10:47 - Z u/l: 0/251 -         2263 - calc.lha
07/20/93 16:21:16 - JR-Comm session finished.

7.5 MagiCall

06/30/93 00:41:46 - MagiCall session started.
06/30/93 00:42:07 - Connected to: E! XS BBS.
06/30/93 00:43:30 - Disconnected.
06/30/93 00:50:55 - Connected to: The Cross.
06/30/93 00:55:16 - Disconnected.
06/30/93 00:55:44 - Connected to: Unknown Carrier.
06/30/93 00:59:23 - Disconnected.
06/30/93 01:00:03 - Connected to: The Cross.
06/30/93 01:04:45 - Disconnected.
06/30/93 01:05:14 - Connected to: AUGS BBS 1.
06/30/93 01:05:37 - Disconnected.
06/30/93 01:06:07 - Connected to: ANet Testbench.
06/30/93 01:08:53 - Disconnected.
06/30/93 01:11:37 - Connected to: The Cross.
06/30/93 01:13:54 - Uploaded file DH1:amanda.
06/30/93 01:14:43 - Disconnected.
06/30/93 01:14:47 - MagiCall session finished.
07/04/93 12:20:45 - PhoneBill LogScan.

7.6 Term action-log

12-jul-93 21:41:02 Start of session (term 3.3 30.4.93).
12-jul-93 21:41:38 Connected to "TUE Dial-in" (040-435049).
12-jul-93 21:47:03 Initiate binary download.
12-jul-93 21:47:11 Initiate binary download.
12-jul-93 21:47:12 Receive file "TERM:/xfer/tinyclock.lha".
12-jul-93 21:47:40 Close file "SPHINX-O-HD:Comm/xfer/tinyclock.lha"
12-jul-93 21:47:53 Initiate binary download.
12-jul-93 21:47:53 Receive file "TERM:/xfer/tinyclock2_0.lha".
12-jul-93 21:48:54 Close file "SPHINX-O-HD:Comm/xfer/tinyclock2_0.lha"
12-jul-93 21:49:05 Hung up the line.
12-jul-93 21:49:10 PhoneBill LogScan.
12-jul-93 21:51:19 End of session.
12-jul-93 21:57:37 Start of session (term 3.3 30.4.93).
12-jul-93 22:24:44 Initiate binary upload.
12-jul-93 22:24:44 Send file "SPHINX-O-HD:Comm/xfer/RECENT".
12-jul-93 22:24:49 Close file "SPHINX-O-HD:Comm/xfer/RECENT".
12-jul-93 22:33:39 Initiate ASCII upload.
12-jul-93 22:33:39 Send file "Workbench:S/user-startup".
12-jul-93 22:33:40 Close file "Workbench:S/user-startup".
12-jul-93 22:46:59 Initiate binary upload.
12-jul-93 23:16:20 Hung up the line.
12-jul-93 23:16:23 End of session.
14-Jul-93 11:05:45 PhoneBill LogScan.

7.7 NComm

Amiga Workbench (05430-24097)
--------------------------------
Login:  Mon Jan 25 21:57:29 1993
Logout: Mon Jan 25 22:08:49 1993
Time online: 00:11:12
Mothership Con. (020-6892708)
--------------------------------
Login:  Tue Jan 26 21:37:13 1993
Logout: Tue Jan 26 21:38:49 1993
Time online: 00:01:36
BBS Waterland(1) (02990-40202)
--------------------------------
Login:  Wed Jan 27 21:10:40 1993
Logout: Wed Jan 27 21:15:41 1993
Time online: 00:05:01
Chat 14400 (5) (046-750624)
--------------------------------
Login:  Thu Jan 28 21:59:57 1993
Logout: Thu Jan 28 22:41:13 1993
Time online: 00:41:16

7.8 Term call-log

Amiga Workbench (05430-24097)
--------------------------------
Login:  Mon Jan 25 21:57:29 1993
Logout: Mon Jan 25 22:08:49 1993
Time online: 00:11:12
Mothership Con. (020-6892708)
--------------------------------
Login:  Tue Jan 26 21:37:13 1993
Logout: Tue Jan 26 21:38:49 1993
Time online: 00:01:36
BBS Waterland(1) (02990-40202)
--------------------------------
Login:  Wed Jan 27 21:10:40 1993
Logout: Wed Jan 27 21:15:41 1993
Time online: 00:05:01
Chat 14400 (5) (046-750624)

--------------------------------
Login:  Thu Jan 28 21:59:57 1993
Logout: Thu Jan 28 22:41:13 1993
Time online: 00:41:16

8 Program updates

Whenever a new release of ‘Phonebill’ becomes available it will be uploaded to:

The Amiga Workbench (SysOp Robert Udo)
+31-(0)5430-24097   (ZyXEL+ 19K2, 24 hours online)
FidoNet:  2:283/410.0
AmigaNet: 39:157/101.0
NLA:      14:105/2.0

I will also try to upload it to the aminet as soon as possible.

9 How to register

As with the previous versions of the program, ‘Phonebill’ v2.0 and higher are shareware. There is one difference: starting from version 2.0 ‘Phonebill’ has been crippled, not all features are available in the unregistered version:

• Binary logfile maintenance is disabled.

• ‘Phonebill’ only calculates the fees with up to two decimals (e.g. $0,15). In the registered version four decimals are supported, which provides a better accuracy.

Many people hate crippled software, including myself. But it just seems that uncrippled software isn’t ‘appreciated’ anymore, because very few people registered. That is why I decided to cripple the program.

Important information: The registration fee for the uncrippled version was $5 US, but I had to raise the price because starting from version 2.0 I will have to send users that register a disk containing a keyfile. Now, starting from 27 September 1993, the registration fee will be $8 US. If I receive any registrations containing $5 US after this date, they will be regarded as a registration for the ‘old’ (pre 2.0) version of the program. So, if you registered with $5 US after this date you will not receive a keyfile: you’ll have to send another $3 US. I would to thank all people that have registered the uncrippled version of the program before this date by sending them a keyfile without having to pay extra.

Registration procedure:

1. Write a note with your complete name and mailing address. If you have access to e-mail (e.g. internet, fidonet, etc.) please specify your e-mail address too. If you want your key to arrive fast I can send it by e-mail, but for security measures you will have to send me your PGP (Pretty Good Privacy) public key first.

2. The registration fee is $8 US, DM 15,-, FL 15,- or the equivalent amount of money in any other

currency. Decide how to transfer the money, e.g.:

• cash.

• a postal money order.

• an EuroCheck (only in FL!), foreign checks are *NOT* accepted!

3. Send this to:

           Raymond Penners
           Berkenlaan 4
           NL-6241 AZ Bunde (The Netherlands)

4. Wait for your keyfile to arrive. You should receive your key within 10 days after the arrival of your registration fee. Please allow up to four weeks, due to unforseen circumstances.

5. Copy the keyfile to the directory where you keep ‘Phonebill’ in or to the ‘L:’ directory. Thanks for registering!

10 Legal stuff

10.1 License

This license applies to the product called ‘Phonebill’, a program for the Amiga computer, published by Raymond Penners under the concepts of ShareWare, and the accompanying documentation, example files and anything else that comes with the original distribution. The terms ‘Program’ and ‘Phonebill’ below, refer to this product. The licensee is addressed as ‘you’.

• You may copy and distribute verbatim copies of the program’s executable code and documentation as you receive it, in any medium, provided that you conspicuously and appropriately publish only the original, unmodified program, with all copyright notices and disclaimers of warranty intact and including all the accompanying documentation, example files and anything else that came with the original.

• You may not copy and/or distribute this program without the accompanying documentation and other additional files that came with the original. You may not copy and/or distribute modified versions of this program.

• You may not copy, modify, sublicense, distribute or transfer the program except as expressly provided under this license. Any attempt otherwise to copy, modify, sublicense, distribute or transfer the program is void, and will automatically terminate your rights to use the program under this license. However, parties who have received copies, or rights to use copies, from you under this license will not have their licenses terminated so long as such parties remain in full compliance.

• By copying, distributing and/or using the program you indicate your acceptance of this license to do so, and all its terms and conditions.

• Each time you redistribute the program, the recipient automatically receives a license from the original licensor to copy, distribute and/or use the program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein.

• You may not disassemble, decompile, re-source or otherwise reverse engineer the program.

• You may use the programs for a period of up to 3 weeks for evaluation. After that, you have to register or quit using the program.

• If you wish to incorporate parts of the programs into other programs, write to the author to ask for permission.

• You agree to cease distributing the program and data involved if requested to do so by the author.

• You may charge a fee to recover distribution costs. The fee for diskette distribution may not be more than the cost to obtain a public domain diskette from Fred Fish.

10.2 No warranty

THERE IS NO WARRANTY FOR THE PROGRAMS, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPY-RIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAMS "AS IS" WITH-OUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PER-FORMANCE OF THE PROGRAMS IS WITH YOU. SHOULD THE PROGRAMS PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRIT-ING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY REDIS-TRIBUTE THE PROGRAMS AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAM-AGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAM-AGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAMS (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAMS TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

10.3 Disclaimer

No warranty, either express or implied, is made with respect to the fitness or merchantability of Phonebill.

Raymond Penners (referred to as ‘the author’), reserves the right to not develop any future versions of Phonebill.

The author will try to make a good faith attempt at correcting any problems if any are discovered, but is in no way required, nor bound to correct them.

The author neither assumes nor accepts any responsibility for the use or misuse of these programs. He will also not be held liable for damages or any compensation beyond the original registration fee due to loss of profit or any other damages arising out of the use, or inability to use this program.

The author will not be liable for any damage arising from the failure of this program to perform as described, or any destruction of other programs or data residing on a system attempting to run the programs. The user of this program uses it at his or her own risk.

11 Acknowledgements

I would like to thank the following people:

• Richard Schiffelers, for being a temporarly P.S.C. :-), replying all the mail, and the 2.04 roms which speed up the development alot! :-)

• Maurice Lahaye, for calling like a mad man, so that I had some logfiles to test my program on.

• Matthias Scheler, for supplying me with the German callrates.

• Philippe Duchein, for the French callrates. (My condolences, the French callrates are really horrible :)

• Johan Billing, for the Swedish callrates

• Henrik Gelius, for sending me information about the MagiCall logfile.

• Jan van den Baard, for his wonderful ‘Gadtoolsbox’.

• Olaf ‘Olsen’ Barthel, for his inspiring ‘Term’.

• Michael Holloway, for sending me his JrComm logfile.

Some quick greetings to: Erwin Penders, Mel Nijsten, Maurice Penners, Walter Konrad, Frans Bouma (Not at ericson anymore? What’s your e-mail adr.?), Robert Udo, Bart Cremers, Arnaud Gouder de Beauregard, Walter Dees, Rino Roufs, Rex (the ONE and ONLY SPHINX-O-Dog), Marcel Meuldijk, Edwin Boesten, Marcel Franquinet ...

Furthermore I would like to say hello to all ‘dudes’ on IRC : Ikke, Bilbo1, Savage, Dense, Milano, Pjotr, ALFred, Mnemo, Xed, SteveVai, Zop, Janne, Outland, Mjoelnir, Juggi, Goldrnr, Lundman, Shocker, Mag, VAG, Herb, Merbot, Mama, and everyone I know but forgot to mention here!

These are all listed in ANALPHABETIC order! I am sorry if I forgot someone (which I’m sure I have) ...

12 Revision history

12.1 Version 1.0

Version 1.0 was released 6-Jun-1993:

• First public release.

12.2 Version 1.07

Version 1.07 was released 14-Jul-1993:

• Compiled on A4000 :-)

• All configuration files now use the IFF format, and are stored in ENV: and ENVARC:.

• The documentation is now in guide format.

• Improved preferences: more options.

• Added requester that warns you on exit if the configuration has been modified.

• Keyboard shortcuts for most of the gadgets.

• Requester warning if <from> date is ’above’ <to> date.

• At the beginning of the report you’ll now see the time period.

• If no calls were found during scanning the report will state an error message. (In V1.00 an empty textfile would be displayed)

• Included built-in rates: German, French, Dutch, Swedish rates.

• Added commandline arguments. Try "Phonebill ?".

• Added menus to the main window.

• Added support for MagiCall logfile.

• In the report produced by "List Calls" sometimes only the phonenumber would be displayed, although a name was provided in the hostmanager.

• If Phonebill was running under a localized workbench, the default from-date ("Yesterday") and to-date ("Today") weren’t accepted anymore.

• Fixed some V39 problems: the busy pointer now appears in the same resolution as the normal mouse pointer, and some problem concerning V39 listviews is fixed too.

• If you were using a bigger system font the rate exception editor would mess up. Also, if you de-zoomed the editor window you wouldn’t see the exception anymore.

12.3 Version 1.08

Version 1.08 was released 1-Aug-1993:

• Added Terminus support.

• Severe bugs were found in in Term, Magicall and NComm scanning (thanks to Benno Zuure, Henrik Gelius and Frank Petzold). I’ve rewritten all log scanning routines to avoid software failures when unknown log messages were encountered.

• The saved configuration files had the wrong version/revision tag. Also, Phonebill would disregard this tag and load any configuration file. So if the configuration files need to be changed in the future, the old version of the program would load the new config files which would probably result in a software failure.

• Under some circumstances phonebill refused to remove a rate, although it wasn’t in use.

12.4 Version 1.1

Version 1.1 was released 16-Aug-1993:

• The host and log lists now show some additional information.

• Added support for Amiga UUCP timelog.

• The pre-defined rate "Sverige: Riks 1" was wrong. The exception "Monday-Friday 18:00-22:00 0.29 per 52 secs" was missing.

12.5 Version 2.0

Version 2.0 was released 27-Sep-1993:

• GUI facelift: windows backfilled, listview doubleclick, changes in the menus.

• Brand new rate editor

• Improved "Do you really want to quit?" requester.

• <From> and <To> date are now saved as strings.

• Iconification.

• Crippled.

• Added JrComm support.

• Binary logfile maintenance.

• Context-sensitive online help, by pressing ‘help’.

• New improved documentation, including ASCII and DVI formats.

• Improved keyboard control: ‘Escape’ will function as a cancel/use action, and ‘return’ as an ok/save action.

• ‘Skip’ in the ‘Unlisted host found’-requester now really skips the host, and not just for once.

• Now ‘Phonebill’ is able to scan a ‘Terminus 2.0b’ logfile correctly (it has some strange ‘LF’ before its connection string).

13 Author Info

Send bug reports, logfiles, suggestions, comments, postcards, flames, contributions, etc. to:

• Snail-Mail:

         Raymond Penners
         Berkenlaan 4
         NL-6241 AZ Bunde (The Netherlands)

• Internet: raymondp@stack.urc.tue.nl

• Fido: 2:283/410.15

• AmigaNet: 39:157/101.15

• NLA: 14:105/201.15

• You may also find me (nickname "Sphinx") on IRC in the #amiga channel.

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1 Upgrading from version 1.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2 Upgrading from version 1.07 - 1.10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

4 Workbench and Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5.1 Project. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.2 Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

6 Control panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

6.1 Main panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 6.2 Preferences panel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 6.3 Log manager panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.4 Log settings panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.5 Host manager panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.6 Host settings panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.7 Rate panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.8 Maintenance panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

7 Supported logfiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

7.1 TrapDoor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 7.2 AmigaUUCP Timelog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 7.3 Terminus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 7.4 JrComm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 7.5 MagiCall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.6 Term action-log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.7 NComm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 7.8 Term call-log. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

8 Program updates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 9 How to register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

10 Legal stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

10.1 License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 10.2 No warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 10.3 Disclaimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

11 Acknowledgements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

12 Revision history. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

12.1 Version 1.0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 12.2 Version 1.07. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 12.3 Version 1.08. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 12.4 Version 1.1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 12.5 Version 2.0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

13 Author Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26